# Protocode: Prototype-Driven Interpretability for Code Generation in LLMs

### v1

@@ 1 Introduction
In recent years, Large Language Models (LLMs) have gained significant traction in the fields of code completion and code filling. This growth has been fueled by the availability of large-scale open-source datasets such as The vault Manh et al. (2023), CodeSearchNet Husain et al. (2020), CodeXGlue Lu et al. (2021) and many others. Alongside these datasets, we have also witnessed the emergence of open-source models designed specifically for code-related tasks, including the CodeLlama series Rozière et al. (2024), Qwen Coder Hui et al. (2024) series, and StarCoder series Li et al. (2023a). In parallel, closed-source models such as GPT-4o OpenAI et al. (2024) and Claude Code Anthropic (2025) have been widely adopted by various big tech companies for generating production-ready code. Despite these advancements, most of these models remain difficult to interpret in the context of code generation. While a variety of interpretability methods have been developed to interpret the outputs generated by LLMs and foster trust in their usage across domains, many of these approaches are generic and not specifically tailored for code generation tasks. Some methods, however, are focused on interpretability in code generation. For instance, Code-Q Palacio et al. (2025) identifies influential tokens that guide the model’s output, but it requires repeated sampling and generation, which introduces significant computational overhead during inference.
@@ 2 Related work
According to Bilal et al. (2025), explainability techniques in AI systems can be broadly divided into three categories: (1) post hoc explanations, (2) intrinsic interpretability, and (3) human-centered explanations. Post hoc explanation methods aim to interpret a model’s decisions after predictions have been made. Common approaches include Local Interpretable Model-Agnostic Explanations (LIME) Ribeiro et al. (2016), Shapley Additive Explanations (SHAP) Lundberg & Lee (2017). LIME provides local explanations by identifying the most important features for a single prediction. Similarly, SHAP evaluates the contribution of each feature by measuring changes in the prediction when features are systematically removed. In addition, gradient-based methods such as SmoothGrad Smilkov et al. (2017) and Integrated Gradients Sundararajan et al. (2017) calculate model gradients with respect to input features to determine the sensitivity of the model’s output to each feature.
@@
A particularly influential method within intrinsic interpretability is Chain-of-Thought (CoT) Wei et al. (2023), which generates intermediate reasoning steps. CoT has been shown to improve both plausibility and task performance compared to demonstrations that provide only the final answers Wei et al. (2023) Cobbe et al. (2021). Building upon this, Self-Consistency Wang et al. (2023) was proposed as an extension of CoT. This method prompts the model to produce multiple reasoning chains and answers, and then selects the final output using a majority vote across the answers. Although effective, Self-Consistency only ensures correctness of the final prediction, without verifying whether the reasoning chains themselves are valid or faithful. To address this, SEA-CoT Wei Jie et al. (2024) was introduced. SEA-CoT evaluates generated reasoning chains based on the implication with the task context and the overlap of the token level, ensuring that both the reasoning process and the final answer align more closely with the task requirements. However, as stated by Jacovi & Goldberg (2020), the reasoning chains from LLM often appear plausible to humans but are not necessarily faithful to the true decision-making process of the LLM. Plausibility refers to how convincing the interpretation is to humans, while faithfulness measures the degree to which it truly represents the internal reasoning of the LLM.
@@ 3 Methodology
To overcome these shortcomings, we turn to the manifold perspective. The manifold hypothesis Cayton (2005) posits that high-dimensional representations lie on low-dimensional manifolds. Leveraging this structure allows prototypes to be sampled from regions that faithfully capture the local geometry of the data, rather than from globally averaged or distorted positions in embedding space. While classical manifold learning techniques such as t-SNE van der Maaten & Hinton (2008), UMAP McInnes et al. (2020), and LLE Roweis & Saul (2000) emphasize neighborhood preservation, they often distort local dependencies or fail to maintain global structure. We therefore adopt a piecewise-linear manifold learning strategy, which decomposes nonlinear manifolds into locally linear regions.
@@ 4 Results
We evaluated the effectiveness of different sampling methods by applying them as in-context learning (ICL) examples on the MBPP test set Austin et al. (2021). To demonstrate the effectiveness of our method we have used 2 sets of models for experimentation, the first set consisting of generic models of Qwen3 Yang et al. (2025), Llama-3.2 AI (2024),Falcon-3 Team (2024) and for the second set we have used code heavy pre-trained models Starcoder-base Li et al. (2023b), Qwen2.5-Coder Hui et al. (2024), Codellama Rozière et al. (2024).
The results are reported on a scale from $[0,1]$, where 0 is the lowest and $1$ is the highest (For instance, 0.1 can be interpreted as 10%). While the numerical margins may appear small at first glance, even modest gains in code completion represent substantial improvements. For context, GPT-4-1106 ope (2023), which is estimated to be at least $1000\times$ larger than the models used for our experiments, achieves a score of $0.786$ on the MBPP test set. This comparison highlights an important distinction: in many benchmarks, partial overlap between a generated solution and the reference solution may yield a nonzero score even if the final answer is incorrect. In contrast, code benchmarks are more stringent, as each generated program is independently evaluated against unit test cases. Therefore, even incremental improvements in $Pass@k$ metrics are highly significant for code generation tasks.
@@ A.1 Manifold learning
Based on the Manifold hypothesis Manifold learning focuses on uncovering low-dimensional structures in high dimensional data. Manifold learning techniques like TSNE van der Maaten & Hinton (2008),UMAP McInnes et al. (2020), LLE Roweis & Saul (2000) and Isomap Tenenbaum et al. (2000) utilize information derived from the linearized neighborhoods of points to construct low dimensional projections of non-linear manifolds in high dimensional data.
The method Piecewise-Linear Manifolds for Deep Metric Learning Bhatnagar & Ahuja (2024) aims to train a neural network to learn a semantic feature space where similar items are close together and dissimilar items are far apart, in an unsupervised manner. This method is based on using linearized neighborhoods of points to construct a piecewise linear manifold, which helps estimate a continuous-valued similarity between data points.

### v2

@@ 1 Introduction
In recent years, Large Language Models (LLMs) have gained significant traction in the fields of code completion and code filling. This growth has been fueled by the availability of large-scale open-source datasets such as The vault Manh et al. (2023), CodeSearchNet Husain et al. (2020), CodeXGlue Lu et al. (2021) and many others. Alongside these datasets, we have also witnessed the emergence of open-source models designed specifically for code-related tasks, including the CodeLlama series rozière2024codellamaopenfoundation, Qwen Coder Hui et al. (2024) series, and StarCoder series Li et al. (2023b). In parallel, closed-source models such as GPT-4o OpenAI et al. (2024) and Claude Code Anthropic (2025) have been widely adopted by various big tech companies for generating production-ready code. Despite these advancements, most of these models remain difficult to interpret in the context of code generation. While a variety of interpretability methods have been developed to interpret the outputs generated by LLMs and foster trust in their usage across domains, many of these approaches are generic and not specifically tailored for code generation tasks. Some methods, however, are focused on interpretability in code generation. For instance, Code-Q Palacio et al. (2025) identifies influential tokens that guide the model’s output, but it requires repeated sampling and generation, which introduces significant computational overhead during inference.
@@ 2 Related work
According to Bilal et al. (2025), explainability techniques in AI systems can be broadly divided into three categories: (1) post hoc explanations, (2) intrinsic interpretability, and (3) human-centered explanations. Post hoc explanation methods aim to interpret a model’s decisions after predictions have been made. Common approaches include Local Interpretable Model-Agnostic Explanations (LIME) Ribeiro et al. (2016), Shapley Additive Explanations (SHAP) Lundberg and Lee (2017). LIME provides local explanations by identifying the most important features for a single prediction. Similarly, SHAP evaluates the contribution of each feature by measuring changes in the prediction when features are systematically removed. In addition, gradient-based methods such as SmoothGrad Smilkov et al. (2017) and Integrated Gradients Sundararajan et al. (2017) calculate model gradients with respect to input features to determine the sensitivity of the model’s output to each feature.
@@
A particularly influential method within intrinsic interpretability is Chain-of-Thought (CoT) Wei et al. (2023), which generates intermediate reasoning steps. CoT has been shown to improve both plausibility and task performance compared to demonstrations that provide only the final answers Wei et al. (2023) Cobbe et al. (2021). Building upon this, Self-Consistency Wang et al. (2023) was proposed as an extension of CoT. This method prompts the model to produce multiple reasoning chains and answers, and then selects the final output using a majority vote across the answers. Although effective, Self-Consistency only ensures correctness of the final prediction, without verifying whether the reasoning chains themselves are valid or faithful. To address this, SEA-CoT Wei Jie et al. (2024) was introduced. SEA-CoT evaluates generated reasoning chains based on the implication with the task context and the overlap of the token level, ensuring that both the reasoning process and the final answer align more closely with the task requirements. However, as stated by Jacovi and Goldberg (2020), the reasoning chains from LLM often appear plausible to humans but are not necessarily faithful to the true decision-making process of the LLM. Plausibility refers to how convincing the interpretation is to humans, while faithfulness measures the degree to which it truly represents the internal reasoning of the LLM.
@@ 3 Methodology
To overcome these shortcomings, we turn to the manifold perspective. The manifold hypothesis Cayton (2005) posits that high-dimensional representations lie on low-dimensional manifolds. Leveraging this structure allows prototypes to be sampled from regions that faithfully capture the local geometry of the data, rather than from globally averaged or distorted positions in embedding space. While classical manifold learning techniques such as t-SNE van der Maaten and Hinton (2008), UMAP McInnes et al. (2020), and LLE Roweis and Saul (2000) emphasize neighborhood preservation, they often distort local dependencies or fail to maintain global structure. We therefore adopt a piecewise-linear manifold learning strategy, which decomposes nonlinear manifolds into locally linear regions.
@@ 4 Results
We evaluated the effectiveness of different sampling methods by applying them as in-context learning (ICL) examples on the MBPP test set Austin et al. (2021). To demonstrate the effectiveness of our method we have used 2 sets of models for experimentation, the first set consisting of generic models of Qwen3 Yang et al. (2025), Llama-3.2 AI (2024),Falcon-3 Team (2024) and for the second set we have used code heavy pre-trained models Starcoder-base Li et al. (2023a), Qwen2.5-Coder Hui et al. (2024), Codellama rozière2024codellamaopenfoundation.
The results are reported on a scale from $[0,1]$, where 0 is the lowest and $1$ is the highest (For instance, 0.1 can be interpreted as 10%). While the numerical margins may appear small at first glance, even modest gains in code completion represent substantial improvements. For context, GPT-4-1106 34, which is estimated to be at least $1000\times$ larger than the models used for our experiments, achieves a score of $0.786$ on the MBPP test set. This comparison highlights an important distinction: in many benchmarks, partial overlap between a generated solution and the reference solution may yield a nonzero score even if the final answer is incorrect. In contrast, code benchmarks are more stringent, as each generated program is independently evaluated against unit test cases. Therefore, even incremental improvements in $Pass@k$ metrics are highly significant for code generation tasks.
@@ 6 Future Works
Acknowledgments
The authors were partially supported by the US National Science Foundation under awards IIS-2520978, GEO/RISE-5239902, the Office of Naval Research Award N00014- 23-1-2007, DOE (ASCR) Award DE-SC0026052, and the DARPA D24AP00325-00. Approved for public release; distribution is unlimited.
@@ A.1 Manifold learning
Based on the Manifold hypothesis Manifold learning focuses on uncovering low-dimensional structures in high dimensional data. Manifold learning techniques like TSNE van der Maaten and Hinton (2008),UMAP McInnes et al. (2020), LLE Roweis and Saul (2000) and Isomap Tenenbaum et al. (2000) utilize information derived from the linearized neighborhoods of points to construct low dimensional projections of non-linear manifolds in high dimensional data.
The method Piecewise-Linear Manifolds for Deep Metric Learning Bhatnagar and Ahuja (2024) aims to train a neural network to learn a semantic feature space where similar items are close together and dissimilar items are far apart, in an unsupervised manner. This method is based on using linearized neighborhoods of points to construct a piecewise linear manifold, which helps estimate a continuous-valued similarity between data points.